OpenStudy (anonymous):

y varies jointly with x and the square of n. When y = 25, x = 3 and n = 4. Find the value of y when x = 5 and n = 10. Round to the nearest whole number. Choose one answer. a. 260 b. 290 c. 320 d. 360

OpenStudy (anonymous):

oh my im so stuck on this question!

OpenStudy (anonymous):

a

OpenStudy (anonymous):

x went up by 5/3 and n2 went up by 100/16. 5/3 * 100/16 * 25 = 260.42

OpenStudy (anonymous):

so...\[y \alpha x ; y \alpha n^2\] therefore the relation is linear in x

OpenStudy (anonymous):

Let K be a constant in $y=k x n^2$. 25 = k 3 16 k = 25/(3 *16) so, \[y = \dfrac{25}{3*16} * 5 * 100\]

OpenStudy (anonymous):

a...using linearity i.e. y/(x*n^2) = const